OpenStudy (anonymous):

Write a quadratic function rule for the given table in problems 1 and 2. 1. (graph below) y=x-6 y=x-4 y=x to the 2nd power - 3 y=x to the 2nd power - 6 2. (graph below) y=x+4 y=x to the 2nd power + 4 y=x to the 2nd power + 6 y=x+6

OpenStudy (anonymous):

1.

OpenStudy (anonymous):

2.

OpenStudy (anonymous):

@greenlegodude57 Help with 2 more?

OpenStudy (anonymous):

@Hotchellerae21

OpenStudy (anonymous):

@kelliegirl33 @manuelshadow sorry guys, but I really need this done

OpenStudy (anonymous):

@pashauna

OpenStudy (anonymous):

@Swiftie.13

OpenStudy (anonymous):

@ShadowLegendX

OpenStudy (anonymous):

@Uri

OpenStudy (anonymous):

@ChiefArnav

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

Have you ever done graphing equations. You know what plugging in any number for x, solving for y, and getting your (x,y) coordinate means ?

OpenStudy (anonymous):

I might not. Can you give an example?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm y=x-6 }\) If x=0, then `y=(0)-6` --> `y=-6` thus `(0,-6)`

OpenStudy (anonymous):

No, I haven't done that.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[y=x-6 \\ y=x-4\] are out, they are not quadratics

OpenStudy (anonymous):

\[y=x^2-3\\ y=x^2-6\]check which one has \((3,3)\)

OpenStudy (anonymous):

y=x^2-6?

OpenStudy (anonymous):

@satellite73